OpenStudy (anonymous):

Perfect square trinomial factoring x^2 + 12x + 36 answer and explain please

jimthompson5910 (jim_thompson5910):

x^2 + 12x + 36 x^2 + 6x+6x + 36 (x^2 + 6x)+(6x + 36) x(x+6)+6(x+6) (x+6)(x+6) (x+6)^2 So x^2 + 12x + 36 factors to (x+6)^2 In other words, x^2 + 12x + 36 = (x+6)^2

OpenStudy (anonymous):

\[(x+a)(x+b)\]where a*b = 36 and a+b = 12 36 is a perfect square. 6^2 = 36. 6+6 = 12. Therefore, this is a perfect square\[(x+6)^2\]

jimthompson5910 (jim_thompson5910):

Note: the numbers 6 and 6 both add to 12 and multiply to 36

OpenStudy (anonymous):

what if the equation was not a perfect square. for example x^2 - 10x +25

jimthompson5910 (jim_thompson5910):

This is a perfect square since x^2-10x+25 = (x-5)^2

OpenStudy (anonymous):

Same approach applies. \[(x+a)(x+b)\]where a*b = 25 and a+b=-10 Here, we can use -5 and -5. These add to -10 and multiply to 25. Again, we have a perfect square.

OpenStudy (anonymous):

so the answer is (x + 5) (x - 15)?

OpenStudy (anonymous):

so the answer is (x + 5) (x - 15)?